OpenStudy (anonymous):

Please help! What mass in grams of Al2O2 (s) will be produced? (Al=26.98u , O=16.00u)

OpenStudy (haseeb96):

=26.98 x 2 + 16 x 2 =53.96+32 =85.96 gm

OpenStudy (haseeb96):

gram is the unit of Gram formula mass whereas a.m.u. is the unit of formula or molecular mass

OpenStudy (anonymous):

See I work it out how you did and i got the same answer! But when i looked at the answer in the back of the book it was 101.96g

OpenStudy (haseeb96):

your question instead of Al2O2 it should be Al2O3 so , = 26.98 x 2 + 16 x 3 = 53.96 + 48 = 101.96 g now it is match with your book answer